# RoadBEV: Road Surface Reconstruction in Bird’s Eye View

### v2

@@ V Experiments
In this section, we comprehensively test performance of the proposed models and verify their superiority for practical RSR applications. We compare RoadBEV-mono with existing monocular depth estimation models, and RoadBEV-stereo with public stereo matching methods. Ablation and comparison studies are implemented to investigate the influence of various parameters.
@@ V-A Implementation Details
We utilize the following metrics to evaluate model performance: absolute error (abs. err.), root mean squared error (RMSE), and ratio of grids with error greater than 0.5 cm (>0.5 cm). We implement the models on PyTorch platform and train on RTX 3090 GPUs. Batch size is 8 for all experiments. Learning rate is set as 8e-4 for RoadBEV-mono, while 5e-4 for RoadBEV-stereo. The optimizer is AdamW with weight decay 1e-4. The OneCycle learning rate scheduler with linear decreasing strategy is used. The images are cropped to 960*528 to meet the size requirements. The RoadBEV-mono is trained for 50 epochs, while 40 for RoadBEV-stereo. Pre-trained weights on ImageNet is loaded. The compared models in Table I are also trained for the same epochs, with their default configurations in the provided codes.
@@ V-B Performance and Comparison
Fig. 9 shows training losses of the proposed two models. They all reach convergence at the set training epochs, corroborating the effectiveness of model structures and loss functions. For the same GT labels and loss functions, the stereo-based model outperforms the monocular-based with lower loss value. The convergence of RoadBEV-stereo is more stable and faster. Comparing with the inexplicable fitting in RoadBEV-mono, the introduction of stereo information significantly contributes more clues to effectively learn road features in vertical direction.
@@
As shown in Table I, we compare with depth estimation and stereo matching methods that ever achieved the SOTA performance on public datasets. Since the compared models finally provide depth in camera’s coordinate, we convert them into BEV and generate elevation maps with the same style as GT labels. For monocular-based road reconstruction, our model promotes the metrics with dominant advantages. The absolute elevation error and RMSE improves by 23.3% and 25.8% than the AdaBins, respectively. The direct elevation estimation in BEV takes effect by uniformly extracting and aggregating features in the vertical direction. The error level at 1.83 cm captures most severe road unevenness that corrupts vehicle ride comfort, but not adequate for slighter road undulations. Moreover, the AdaBins achieves higher accuracy than the others using direct regression, verifying the necessity of classification on bins.
@@
Method Abs. err. (cm) RMSE (cm) >0.5 cm (%)
Mono LapDepth[45] 2.81 3.12 85.3
PixelFormer[46] 2.65 2.86 82.0
iDisc[47] 2.64 2.88 84.3
AdaBins[44] 2.59 2.79 82.4
RoadBEV(Ours) 1.83 2.07 78.6
Stereo IGEV-Stereo[48] 0.651 0.797 49.5
PSMNet[49] 0.654 0.785 50.1
CFNet[50] 0.647 0.760 50.8
ACVNet[51] 0.596 0.723 46.2
GwcNet[52] 0.588 0.711 44.9
DVANet[13] 0.546 0.685 40.9
RoadBEV (Ours) 0.503 0.609 37.0
@@ V-C Visualization of Road Reconstruction
The reconstructed road elevation maps by RoadBEV-mono are visualized in Fig. 11. Road surface structures and trends are accurately captured without holistic bias. Inferred map for the first sample, representing plane road surface without significant unevenness, shows stable elevations. This is crucial for downstream tasks as false detections may leads to unnecessary actions. The elongate crack in the second sample, as well as the abrupt bump at far distance, are recovered with high accuracy, as the residual map does not show prominent errors in the corresponding areas. The rut in the third sample is also caught by our model, verifying its capability in recovering both global structures and local fine variations. The straight speed bump in the forth sample coincides well with GT. Although the amplitude error is relatively big in the middle part as indicated by the residual map, it can still be clearly identified as a bump. The maximum absolute error can be bounded in 2.5 cm. Post-processing methods can be implemented to refine the bump’s profile.
@@
Fig. 12 shows the reconstruction results by RoadBEV-stereo. The second column shows the point clouds with color in BEV, which are obtained by stereo matching model GwcNet. The last column visualizes reconstructed 3D road meshes. Comparing with those in Fig. 11, the recovered elevation maps are smoother without cluttered noise or unexpected patterns. The RoadBEV-stereo is capable of preserving more detailed structures than RoadBEV-mono. The shapes and edges of potholes in the second and last samples are precisely recovered. As illustrated by the 3D meshes, the slopes from pothole edges to centers are also captured with exactness. In the third sample, the sharp steep between the cracked and even areas are clearly identifiable. All the presented samples demonstrate the capability of our method in handling complicated road patterns.
@@ VI Limitations and Prospects
Visualizations in Fig. 10 indicate that the increasing trend of error w.r.t. longitudinal distance still exists for both the monocular and stereo-based. This is an inherent drawback of perspective camera, where texture and structure details are preserved at near distance while lost at far distance. Although the BEV paradigm directly reconstructs road surface from a top-down view, the features are still extracted from perspective images. BEV is a promising way to suppress this phenomenon, with further efforts and more advanced strategies.
In our implementations, we only adopt frames at current time. RoadBEV-stereo demonstrates that introducing information from other perspectives profoundly promotes model performance. Therefore, utilizing sequence images is expected to bring further enhancements. In view transformation, we project voxel centers onto image plane and index the corresponding pixel features. For more accurate feature query, projecting multiple voxel locations and fusing nearby pixel features deserve attention.
We investigate model performance at the same horizontal resolution (i.e., 164*64 grids). Like stereo matching, the BEV volume can be first constructed at lower resolution and then integrally interpolated to full resolution, which effectively reduces computation. In this paper, we focus only on reconstructing road geometry structure, i.e., elevation. For future research, joint geometry and texture reconstruction can be explored, with recent technologies like NeRF [53] and 3D Gaussian Splatting [54].
Data scale and diversity are the most essential problems to be considered for real-scenario applications. We utilize our previous work, RSRD, to produce this algorithm and application prototype. Although typical road conditions are covered, the diversity is still inadequate, especially various corner cases. Efforts should be put to contribute more high-quality road surface data.
@@ VII Conclusion
In this paper, we reconstruct road surface elevation in Bird’s Eye View for the first time. Two models named RoadBEV-mono and RoadBEV-stereo based on monocular and stereo images are proposed and analyzed, respectively. We reveal that monocular estimation and stereo matching in BEV have the same mechanism with that in perspective view, while improves by narrowing search range and digging features directly in elevation direction. Comprehensive experiments on real-world dataset validate the feasibility and superiority of proposed, BEV volume, estimation head, and parameter settings. For monocular camera, the reconstruction performance in BEV get 50% promotion than that in perspective view. Meanwhile, in BEV, the performance with stereo camera is three times of the monocular. Insightful analysis and instructions about the models are provided. Our pioneering exploration also extends valuable references to further research and applications associated with BEV perception, 3D reconstruction and 3D detection.

### v3

@@ V Experiments
In this section, we comprehensively test performance of the proposed models and verify their superiority for practical RSR applications. We compare RoadBEV-mono with existing monocular depth estimation models, and RoadBEV-stereo with stereo matching methods. Ablation and comparison studies are implemented to investigate the influence of various parameters.
@@ V-A Implementation Details
We utilize the following metrics to evaluate model performance: absolute error (abs. err.), root mean squared error (RMSE), ratio of grids with error greater than 0.5 cm (>0.5 cm), and frames per second (FPS) in inference. We implement the models on PyTorch platform and train on RTX 3090 GPUs. Batch size is 8 for all experiments. Learning rate is set as 8e-4 for RoadBEV-mono, while 5e-4 for RoadBEV-stereo. The optimizer is AdamW with weight decay 1e-4. The OneCycle learning rate scheduler with linear decreasing strategy is used. The images are cropped to 960*528 to meet the size requirements. The RoadBEV-mono is trained for 50 epochs, while 40 for RoadBEV-stereo. Pre-trained weights on ImageNet is loaded. The compared models in Table I are also trained for the same epochs, with their default configurations in the provided codes.
@@ V-B Performance and Comparison
Fig. 9 shows training losses of the proposed two models. They all reach convergence at the set training epochs, corroborating the effectiveness of model structures and loss functions. For the same GT labels and loss functions, the stereo-based model outperforms the monocular-based with lower loss value. The convergence of RoadBEV-stereo is more stable and faster. Comparing with the direct but inexplicable fitting in RoadBEV-mono, the introduction of stereo information significantly contributes more clues to learn road features in vertical direction.
@@
As shown in Table I, we compare with depth estimation and stereo matching methods that ever achieved the SOTA performance on public datasets. Since the compared models finally provide depth in camera’s coordinate, we convert them into BEV and generate elevation maps with the same style as GT labels. For monocular-based road reconstruction, our model promotes the metrics with dominant advantages. The absolute elevation error and RMSE reduces by 29.3% and 25.8% than the AdaBins, respectively. The direct elevation estimation in BEV takes effect by uniformly extracting and aggregating features in the vertical direction. The error level at 1.83 cm captures most severe road unevenness that corrupts vehicle ride comfort, but not adequate for slighter road undulations. Moreover, the AdaBins achieves higher accuracy than the others using direct regression, verifying the necessity of classification on bins.
@@
Method Abs. err. (cm) RMSE (cm) >0.5 cm (%) FPS
Mono LapDepth[53] 2.81 3.12 85.3 83.2
PixelFormer[54] 2.65 2.86 82.0 43.0
iDisc[55] 2.64 2.88 84.3 12.3
AdaBins[52] 2.59 2.79 82.4 21.5
RoadBEV(Ours) 1.83 2.07 78.6 26.8
Stereo IGEV-Stereo[56] 0.651 0.797 49.5 4.6
PSMNet[57] 0.654 0.785 50.1 11.6
CFNet[58] 0.647 0.760 50.8 6.8
ACVNet[59] 0.596 0.723 46.2 12.0
GwcNet[60] 0.588 0.711 44.9 15.6
DVANet[17] 0.546 0.685 40.9 8.7
RoadBEV (Ours) 0.503 0.609 37.0 8.0
@@
As for inference, our RoadBEV-mono achieves 26.8 FPS, which is promising for real-vehicle applications. The inference speed of RoadBEV-stereo is low due to the 3D convolutions requiring much computation. The inference speed is expected to be promoted in embedded deployment with special operation optimization. Despite that, since the camera has preview distance at about 7 m, the relative low inference speed is thereby compensated. Road elevation is still continuously available for the downstream planning and control units.
@@ V-C Visualization of Road Reconstruction
The reconstructed road elevation maps by RoadBEV-mono are visualized in Fig. 11. Road surface structures and trends are accurately captured without holistic bias. Inferred map for the first sample, representing plane road surface without significant unevenness, shows stable elevations. This is crucial for downstream tasks as false detections may leads to unnecessary actions. The elongate crack in the second sample, as well as the abrupt bump at far distance, are recovered with high accuracy, as the residual map does not show prominent errors in the corresponding areas. The rut in the third sample is also caught, verifying its capability in recovering both global structures and local fine variations. The straight speed bump in the forth sample coincides well with GT. Although the amplitude error is relatively big in the middle part, it can still be clearly identified as a bump. The maximum absolute error can be bounded in 2.5 cm. Post-processing methods can be implemented to refine the bump’s profile.
@@
Fig. 12 shows the reconstruction results by RoadBEV-stereo. The second column shows the point clouds with color in BEV, which are obtained by GwcNet [60]. The last column visualizes reconstructed 3D road meshes. Comparing with those in Fig. 11, the recovered elevation maps are smoother without cluttered noise or unexpected patterns. The RoadBEV-stereo is capable of preserving more detailed structures than RoadBEV-mono. The shapes and edges of potholes in the second and last samples are precisely recovered. As illustrated by the 3D meshes, the slopes from pothole edges to centers are also captured with exactness. In the third sample, the sharp steep between the cracked and even areas are clearly identifiable. All the presented samples demonstrate the capability of our method in handling complicated road patterns.
@@ VI Limitations and Prospects
Visualizations in Fig. 10 indicate that the increasing trend of error w.r.t. longitudinal distance still exists for both the monocular and stereo-based. This is an inherent drawback of perspective camera, where texture and structure details are lost at far distance. Although the BEV paradigm directly reconstructs road surface from a top-down view, the features are still extracted from perspective images. BEV is a promising way to suppress this phenomenon, with further efforts and more advanced strategies.
In our implementations, we only adopt frames at current time. RoadBEV-stereo demonstrates that introducing other perspectives profoundly promotes model performance. Therefore, utilizing sequence images is expected to bring further enhancements. In view transformation, we project voxel centers onto image plane and index the corresponding pixel features. For more accurate feature query, fusing nearby pixel features with spatial cross attention deserve exploration [42].
We investigate model performance at the same horizontal resolution (i.e., 164*64 grids). Similar to stereo matching, the BEV volume can be first constructed at lower resolution and then integrally interpolated to full resolution, which effectively reduces computation. In this paper, we focus only on reconstructing road geometry structure, i.e., elevation. For future research, joint geometry and texture reconstruction can be explored, with recent technologies like NeRF [61, 62, 63] and 3D Gaussian Splatting [64, 65].
We utilize our previous work, RSRD, to produce this algorithm and application prototype. Although typical road conditions are covered, the diversity is still inadequate, especially various corner cases. Efforts should be put to contribute more high-quality road surface data.
@@ VII Conclusion
In this paper, we reconstruct road surface elevation in Bird’s Eye View for the first time. Two models named RoadBEV-mono and RoadBEV-stereo based on monocular and stereo images are proposed and analyzed, respectively. We reveal that monocular estimation and stereo matching in BEV have the same mechanism with that in perspective view, while improves by narrowing search range and digging features directly in elevation direction. Comprehensive experiments on real-world dataset validate the feasibility and superiority of proposed, BEV volume, estimation head, and parameter settings. For monocular camera, the reconstruction error in BEV reduces by about 30% than that in perspective view. Meanwhile, in BEV, the accuracy performance with stereo camera is more than three times of the monocular. Insightful analysis and instructions about the models are provided. Our pioneering exploration also extends valuable references to further research and applications associated with BEV perception, 3D reconstruction and 3D detection.